# MedPT: A Massive Medical Question Answering Dataset for Brazilian-Portuguese Speakers

### v1

@@ 5. Experiments
To test MedPT dataset under downstream applications, we formulate a multi-class classification task designed to predict the appropriate medical subspecialty from a patient’s question. This task simulates a critical real-world use case: the automated routing of patient inquiries to the correct clinical department. To evaluate model robustness against varying levels of label complexity, we create three experimental tiers based on the Top 5, Top 10, and Top 20 most frequent subspecialties in the dataset.
@@ 5.1. Experimental Setup
For all experiments, we use the Qwen3-1.7B Yang et al. (2025) due to its balance between computational efficiency and performance, offering competitive results while maintaining a substantially lower training and inference cost compared to larger-scale models. We then benchmark its performance across three distinct methodological settings corresponding to the Top 5, Top 10, and Top 20 most frequent subspecialties. Each split as divided into an 80% training set and a 20% held-out test set. The evaluation settings are as follows:
@@
Few-Shot (In-Context Learning): The base model is prompted with instructions along with a small number of dynamically-sampled examples from the training set. We evaluate its in-context learning capability with 1, 3, and 5 examples to measure performance with varying levels of contextual information.
@@ 5.2. Results
The experimental results, summarized in Table 3, reveal a clear and consistent performance hierarchy across all tested configurations. Fine-tuning the Qwen3-1.7B model on the MedPT dataset yields the highest performance by a significant margin, establishing a strong benchmark and demonstrating the dataset’s value for model specialization. As expected, performance for all methods degrades gracefully as the task difficulty increases from 5 to 20 classes. Nonetheless, the fine-tuned model maintains an exceptional 0.94 F1-score on the most challenging 20-class task, underscoring the quality of the learned representations.
@@ 6. Discussion
The experimental results confirm the high utility of MedPT for developing specialized medical language models. The exceptional performance of the fine-tuned model, achieving a 0.94 F1-score on the challenging 20-class task, directly validates the dataset’s quality. This demonstrates that the linguistic patterns and class labels within MedPT are consistent and highly learnable, providing a robust foundation for domain-specific classifiers. Furthermore, our findings offer practical implications for deployment: the substantial performance gains from in-context learning confirm MedPT’s value for resource-light applications, while the observed diminishing returns between 3-shot and 5-shot prompts suggest an optimal efficiency-performance trade-off.
A deeper qualitative analysis of the fine-tuned model’s errors revealed that misclassifications were not random but systematic, occurring between semantically and clinically related subspecialties. We observed notable confusion between conditions with high comorbidity (e.g., Anxiety and Depression), anatomical proximity (e.g., Umbilical Hernia and Inguinal Hernia), or overlapping symptoms (e.g., Gastritis and Gastroesophageal Reflux). This pattern indicates that the model is capturing a meaningful semantic space of medical concepts, where its errors reflect genuine clinical ambiguity. This semantic richness underscores MedPT’s potential to support future research into more complex tasks, such as fine-grained medical entity disambiguation.
@@ 7. Conclusion
This high utility was then confirmed through multi-faceted validation. Quantitatively, downstream classification experiments established a strong benchmark, with models fine-tuned on MedPT achieving an outstanding 0.94 F1-score on a 20-class task. This, combined with strong few-shot ICL results, proves the dataset’s practical value. By publicly releasing MedPT, we provide a foundational resource to the research community. We aim to foster the development of more equitable, accurate, and safe medical NLP technologies, enabling a new generation of models that can understand the specific clinical and linguistic nuances of the Portuguese-speaking world.
@@ 8. Limitations
We acknowledge several limitations in this work that also present clear avenues for future research. First, while our experimental validation focused exclusively on a multi-class classification task demonstrates the dataset’s utility for a crucial application like inquiry routing, its performance on complex generative tasks, such as medical question-answering and patient-doctor dialogue, remains to be explored.
Second, regarding data quality, although all answers originate from verified physicians, a secondary, independent clinical review of the responses was not conducted. Such a re-validation, even on a representative subset of the data, could further bolster the dataset’s reliability and ensure the clinical accuracy of the provided information.
Finally, our deduplication process was based on removing exact-match samples or token-count strategy. Future iterations of this work could employ more sophisticated near-duplication detection techniques. While this would likely reduce the dataset’s total size, it could also enhance its cross-sample linguistic diversity by filtering out highly similar, redundant entries.
@@ 10. Bibliographical References
- Abdelhay et al. (2023) Mohammed Abdelhay, Ammar Mohammed, and Hesham A Hefny. 2023. Deep learning for arabic healthcare: Medicalbot. Social Network Analysis and Mining, 13(1):71.
- Achiam et al. (2023) Josh Achiam, Steven Adler, Sandhini Agarwal, Lama Ahmad, Ilge Akkaya, Florencia Leoni Aleman, Diogo Almeida, Janko Altenschmidt, Sam Altman, Shyamal Anadkat, et al. 2023. Gpt-4 technical report. arXiv preprint arXiv:2303.08774.
- Agarwal et al. (2025) Sandhini Agarwal, Lama Ahmad, Jason Ai, Sam Altman, Andy Applebaum, Edwin Arbus, Rahul K Arora, Yu Bai, Bowen Baker, Haiming Bao, et al. 2025. gpt-oss-120b & gpt-oss-20b model card. arXiv preprint arXiv:2508.10925.
@@
- Alhuzali et al. (2024) Hassan Alhuzali, Ashwag Alasmari, and Hamad Alsaleh. 2024. Mentalqa: An annotated arabic corpus for questions and answers of mental healthcare. IEEE Access.
@@
- Alkouz et al. (2019) Balsam Alkouz, Zaher Al Aghbari, and Jemal Hussien Abawajy. 2019. Tweetluenza: Predicting flu trends from twitter data. Big Data Mining and Analytics, 2(4):273–287.
@@
- Cascella et al. (2024) Marco Cascella, Federico Semeraro, Jonathan Montomoli, Valentina Bellini, Ornella Piazza, and Elena Bignami. 2024. The breakthrough of large language models release for medical applications: 1-year timeline and perspectives. Journal of Medical Systems, 48(1):22.
@@
- Consoli et al. (2022) Bernardo Scapini Consoli, Henrique Dias P dos Santos, Ana Helena DPS Ulbrich, Renata Vieira, and Rafael Heitor Bordini. 2022. Brateca (brazilian tertiary care dataset): a clinical information dataset for the portuguese language. In Proceedings of the 13th Conference on Language Resources and Evaluation (LREC 2022), Marseille, 20-25 June, 2022, França.
@@
- Cury et al. (2021) Ricardo C Cury, Istvan Megyeri, Tony Lindsey, Robson Macedo, Juan Batlle, Shwan Kim, Brian Baker, Robert Harris, and Reese H Clark. 2021. Natural language processing and machine learning for detection of respiratory illness by chest ct imaging and tracking of covid-19 pandemic in the united states. Radiology: Cardiothoracic Imaging, 3(1):e200596.
@@
- da Costa Lopes (2019) Fábio André da Costa Lopes. 2019. Contributions to clinical information extraction in portuguese: Corpora, named entity recognition, word embeddings. Master’s thesis, Universidade de Coimbra (Portugal).
@@
- Enshaei and Naderkhani (2024) Nastaran Enshaei and Farnoosh Naderkhani. 2024. The role of data quality for reliable ai performance in medical applications. IEEE Reliability Magazine, 1(3):24–28.
@@
- Färber et al. (2025) Fernanda B Färber, Julia S Dollis, Pedro SFB Ribeiro, Iago A Brito, Rafael T Sousa, and Arlindo R Galvão Filho. 2025. Iaramed: A women’s healthcare chatbot for portuguese speakers. In Simpósio Brasileiro de Computação Aplicada à Saúde (SBCAS), pages 931–942. SBC.
@@
- Ghassabi et al. (2025) Mehrdad Ghassabi, Pedram Rostami, Hamidreza Baradaran Kashani, Amirhossein Poursina, Zahra Kazemi, and Milad Tavakoli. 2025. Leveraging online data to enhance medical knowledge in a small persian language model. arXiv preprint arXiv:2505.16000.
@@
- Guo et al. (2018) Haihong Guo, Xu Na, and Jiao Li. 2018. Qcorp: an annotated classification corpus of chinese health questions. BMC medical informatics and decision making, 18(Suppl 1):16.
@@
- He et al. (2019) Junqing He, Mingming Fu, and Manshu Tu. 2019. Applying deep matching networks to chinese medical question answering: a study and a dataset. BMC medical informatics and decision making, 19(Suppl 2):52.
- He et al. (2020) Xuehai He, Shu Chen, Zeqian Ju, Xiangyu Dong, Hongchao Fang, Sicheng Wang, Yue Yang, Jiaqi Zeng, Ruisi Zhang, Ruoyu Zhang, et al. 2020. Meddialog: Two large-scale medical dialogue datasets. arXiv preprint arXiv:2004.03329.
- Ji et al. (2019) Bin Ji, Rui Liu, Shasha Li, Jie Yu, Qingbo Wu, Yusong Tan, and Jiaju Wu. 2019. A hybrid approach for named entity recognition in chinese electronic medical record. BMC medical informatics and decision making, 19(Suppl 2):64.
- Jin et al. (2019) Qiao Jin, Bhuwan Dhingra, Zhengping Liu, William W Cohen, and Xinghua Lu. 2019. Pubmedqa: A dataset for biomedical research question answering. arXiv preprint arXiv:1909.06146.
- Johnson et al. (2023) Alistair EW Johnson, Lucas Bulgarelli, Lu Shen, Alvin Gayles, Ayad Shammout, Steven Horng, Tom J Pollard, Sicheng Hao, Benjamin Moody, Brian Gow, et al. 2023. Mimic-iv, a freely accessible electronic health record dataset. Scientific data, 10(1):1.
- Khan et al. (2021) Samir Khan, Seiji Tsutsumi, Takehisa Yairi, and Shinichi Nakasuka. 2021. Robustness of ai-based prognostic and systems health management. Annual Reviews in Control, 51:130–152.
- Liu et al. (2025) Qiming Liu, Ruirong Yang, Qin Gao, Tengxiao Liang, Xiuyuan Wang, Shiju Li, Bingyin Lei, and Kaiye Gao. 2025. A review of applying large language models in healthcare. IEEE Access, 13:6878–6892.
- Luo et al. (2022) Ling Luo, Po-Ting Lai, Chih-Hsuan Wei, Cecilia N Arighi, and Zhiyong Lu. 2022. Biored: a rich biomedical relation extraction dataset. Briefings in Bioinformatics, 23(5):bbac282.
- Magna et al. (2020) Andrés Alejandro Ramos Magna, Héctor Allende-Cid, Carla Taramasco, Carlos Becerra, and Rosa L Figueroa. 2020. Application of machine learning and word embeddings in the classification of cancer diagnosis using patient anamnesis. Ieee Access, 8:106198–106213.
- Mulyar et al. (2021) Andriy Mulyar, Ozlem Uzuner, and Bridget McInnes. 2021. Mt-clinical bert: scaling clinical information extraction with multitask learning. Journal of the American Medical Informatics Association, 28(10):2108–2115.
- Naveen and Trojovskỳ (2024) Palanichamy Naveen and Pavel Trojovskỳ. 2024. Overview and challenges of machine translation for contextually appropriate translations.” iscience 27 (10): 110878.
- Nazi and Peng (2024) Z Al Nazi and W Peng. 2024. Large language models in healthcare and medical domain: A review. informatics, 11, 57.
- Nguyen et al. (2022) Nhung Thi-Hong Nguyen, Phuong Phan-Dieu Ha, Luan Thanh Nguyen, Kiet Van Nguyen, and Ngan Luu-Thuy Nguyen. 2022. Spbertqa: A two-stage question answering system based on sentence transformers for medical texts. In International Conference on Knowledge Science, Engineering and Management, pages 371–382. Springer.
- Nguyen et al. (2023) Vincent Nguyen, Sarvnaz Karimi, Maciej Rybinski, and Zhenchang Xing. 2023. Medredqa for medical consumer question answering: Dataset, tasks, and neural baselines. In Proceedings of the 13th International Joint Conference on Natural Language Processing and the 3rd Conference of the Asia-Pacific Chapter of the Association for Computational Linguistics (Volume 1: Long Papers), pages 629–648.
- Oliveira et al. (2022) Lucas Emanuel Silva e Oliveira, Ana Carolina Peters, Adalniza Moura Pucca Da Silva, Caroline Pilatti Gebeluca, Yohan Bonescki Gumiel, Lilian Mie Mukai Cintho, Deborah Ribeiro Carvalho, Sadid Al Hasan, and Claudia Maria Cabral Moro. 2022. Semclinbr-a multi-institutional and multi-specialty semantically annotated corpus for portuguese clinical nlp tasks. Journal of Biomedical Semantics, 13(1):13.
- Paiola et al. (2024) Pedro Henrique Paiola, Gabriel Lino Garcia, Joao Renato Ribeiro Manesco, Mateus Roder, Douglas Rodrigues, and João Paulo Papa. 2024. Adapting llms for the medical domain in portuguese: A study on fine-tuning and model evaluation. arXiv preprint arXiv:2410.00163.
- Pal et al. (2022) Ankit Pal, Logesh Kumar Umapathi, and Malaikannan Sankarasubbu. 2022. Medmcqa: A large-scale multi-subject multi-choice dataset for medical domain question answering. In Proceedings of the Conference on Health, Inference, and Learning, volume 174 of Proceedings of Machine Learning Research, pages 248–260. PMLR.
- Preston et al. (2023) Sam Preston, Mu Wei, Rajesh Rao, Robert Tinn, Naoto Usuyama, Michael Lucas, Yu Gu, Roshanthi Weerasinghe, Soohee Lee, Brian Piening, et al. 2023. Toward structuring real-world data: Deep learning for extracting oncology information from clinical text with patient-level supervision. Patterns, 4(4).
- Ryan Conmy et al. (2023) Philippa Ryan Conmy, Berk Ozturk, Tom Lawton, and Ibrahim Habli. 2023. The impact ofătraining data shortfalls onăsafety ofăai-based clinical decision support systems. In Computer Safety, Reliability, and Security, pages 213–226, Cham. Springer Nature Switzerland.
- Schwabe et al. (2024) Daniel Schwabe, Katinka Becker, Martin Seyferth, Andreas Klaß, and Tobias Schaeffter. 2024. The metric-framework for assessing data quality for trustworthy ai in medicine: a systematic review. NPJ digital medicine, 7(1):203.
- Singhal et al. (2022) Karan Singhal, Shekoofeh Azizi, Tao Tu, S Sara Mahdavi, Jason Wei, Hyung Won Chung, Nathan Scales, Ajay Tanwani, Heather Cole-Lewis, Stephen Pfohl, et al. 2022. Large language models encode clinical knowledge. arXiv preprint arXiv:2212.13138.
- Singhal et al. (2025) Karan Singhal, Tao Tu, Juraj Gottweis, Rory Sayres, Ellery Wulczyn, Mohamed Amin, Le Hou, Kevin Clark, Stephen R Pfohl, Heather Cole-Lewis, et al. 2025. Toward expert-level medical question answering with large language models. Nature Medicine, 31(3):943–950.
- Wang et al. (2020) Jing Wang, Huan Deng, Bangtao Liu, Anbin Hu, Jun Liang, Lingye Fan, Xu Zheng, Tong Wang, and Jianbo Lei. 2020. Systematic evaluation of research progress on natural language processing in medicine over the past 20 years: bibliometric study on pubmed. Journal of medical Internet research, 22(1):e16816.
- Wu et al. (2019) Stephen Wu, Kirk Roberts, Surabhi Datta, Jingcheng Du, Zongcheng Ji, Yuqi Si, Sarvesh Soni, Qiong Wang, Qiang Wei, Yang Xiang, Bo Zhao, and Hua Xu. 2019. Deep learning in clinical natural language processing: a methodical review. Journal of the American Medical Informatics Association, 27(3):457–470.
- Yang et al. (2025) An Yang, Anfeng Li, Baosong Yang, Beichen Zhang, Binyuan Hui, Bo Zheng, Bowen Yu, Chang Gao, Chengen Huang, Chenxu Lv, et al. 2025. Qwen3 technical report. arXiv preprint arXiv:2505.09388.
- Zhang et al. (2018) Sheng Zhang, Xin Zhang, Hui Wang, Lixiang Guo, and Shanshan Liu. 2018. Multi-scale attentive interaction networks for chinese medical question answer selection. IEEE Access, 6:74061–74071.
- Zhu et al. (2023) Enwei Zhu, Qilin Sheng, Huanwan Yang, Yiyang Liu, Ting Cai, and Jinpeng Li. 2023. A unified framework of medical information annotation and extraction for chinese clinical text. Artificial Intelligence in Medicine, 142:102573.
@@
- İncidelen and Aydoğan (2024) Mert İncidelen and Murat Aydoğan. 2024. Developing question-answering models in low-resource languages: A case study on turkish medical texts using transformer-based approaches. In 2024 8th International Artificial Intelligence and Data Processing Symposium (IDAP), pages 1–4.

### v2

@@ 5. Experiments
To test MedPT dataset under downstream applications, we formulate a multi-class classification task designed to predict the appropriate medical specialty from a patient’s question. This task simulates a critical real-world use case: the automated routing of patient inquiries to the correct clinical department. To evaluate model robustness against varying levels of label complexity, we create three experimental tiers based on the Top 5, Top 10, and Top 20 most frequent subspecialties in the dataset.
@@ 5.1. Experimental Setup
For all experiments, we use the Qwen3 1.7B Yang et al. (2025) due to its balance between computational efficiency and performance, offering competitive results while maintaining a substantially lower training and inference cost compared to larger-scale models. We then benchmark its performance across three distinct methodological settings corresponding to the Top 5, Top 10, and Top 20 most frequent subspecialties. Each split as divided into an 80% training set and a 20% held-out test set. The evaluation settings are as follows:
@@
Few-Shot (In-Context Learning): The base model is prompted with instructions along with a small number of dynamically-sampled examples from the training set. We evaluate its in-context learning capability with 3 and 5 examples to measure performance with varying levels of contextual information.
@@ 5.2. Results
The experimental results, summarized in Table 3, reveal a clear and consistent performance hierarchy across all tested configurations. Fine-tuning the Qwen3-1.7B model on the MedPT dataset yields the highest performance by a significant margin, establishing a strong benchmark and demonstrating the dataset’s value for model specialization. As expected, performance for all methods degrades systematically as the task difficulty increases from 5 to 20 classes. Nonetheless, the fine-tuned model maintains an exceptional 0.94 F1-score on the most challenging 20-class task, underscoring the quality of the learned representations.
@@ 6. Discussion
The experimental results confirm the high utility of MedPT for developing specialized medical language models. The exceptional performance of the fine-tuned model, achieving an F1-score of 0.94 on the medical specialty classification task involving 20 specialties, directly validates the dataset’s quality. This demonstrates that the linguistic patterns and class labels within MedPT are consistent and highly learnable, providing a robust foundation for domain-specific classifiers. Furthermore, our findings offer practical implications for deployment: the substantial performance gains from in-context learning confirm MedPT’s value for resource-light applications, while the observed diminishing returns between 3-shot and 5-shot prompts suggest an optimal efficiency-performance trade-off.
A deeper qualitative analysis of the fine-tuned model’s errors revealed that misclassifications were not random but systematic, occurring between semantically and clinically related subspecialties. We observed notable confusion between conditions with high comorbidity (e.g., Anxiety and Depression), anatomical proximity (e.g., Umbilical Hernia and Inguinal Hernia), or overlapping symptoms (e.g., Gastritis and Gastroesophageal Reflux). This pattern indicates that the model is capturing a meaningful semantic space of medical concepts, where its errors reflect genuine clinical ambiguity, underscoring MedPT’s potential to support future research into more complex tasks, such as fine-grained medical entity disambiguation.
@@ 7. Conclusion
This high utility was then confirmed through multi-faceted validation. In downstream classification experiments, a 1.7B-parameter model fine-tuned on MedPT achieved an F1-score of 0.94 on the medical specialty classification task across 20 specialties, compared to 0.60 for the base model. This, combined with strong few-shot ICL results, proves the dataset’s practical value. By publicly releasing MedPT, we provide a foundational resource to the research community. We aim to support the development of more equitable, accurate, and safe medical NLP technologies, enabling a new generation of models that can understand the specific clinical and linguistic nuances of the Portuguese-speaking world.
@@ 8. Limitations
We acknowledge several limitations in this work that also present clear avenues for future research. First, while our experimental validation focused exclusively on a multi-class classification task demonstrates the dataset’s utility for a crucial application like inquiry routing, its performance on generative tasks, such as medical question-answering and patient-doctor dialogue, remains to be explored. Second, our deduplication process was based on removing exact-match samples instead of more sophisticated near-duplication detection techniques. Finally, regarding data quality, although all answers originate from verified physicians, a secondary, independent clinical review of the responses could further reinforce the dataset’s reliability and ensure the clinical accuracy of the provided information.
@@ 9. Ethical Considerations
10. Acknowledgments
This work has been fully/partially funded by the project Research and Development of Algorithms for Construction of Digital Human Technological Components supported by Advanced Knowledge Center in Immersive Technologies (AKCIT), with financial resources from the PPI IoT of the MCTI grant number 057/2023, signed with EMBRAPII.
@@ 11. Bibliographical References
- Deep learning for arabic healthcare: medicalbot. Social Network Analysis and Mining 13 (1), pp. 71. Cited by: §1, §2.
- Gpt-4 technical report. arXiv preprint arXiv:2303.08774. Cited by: §3.
- Gpt-oss-120b & gpt-oss-20b model card. arXiv preprint arXiv:2508.10925. Cited by: §3.3.
@@
- Mentalqa: an annotated arabic corpus for questions and answers of mental healthcare. IEEE Access. Cited by: §3.3.
@@
- Tweetluenza: predicting flu trends from twitter data. Big Data Mining and Analytics 2 (4), pp. 273–287. External Links: Document Cited by: §2.
@@
- The breakthrough of large language models release for medical applications: 1-year timeline and perspectives. Journal of Medical Systems 48 (1), pp. 22. Cited by: §2.
@@
- Brateca (brazilian tertiary care dataset): a clinical information dataset for the portuguese language. In Proceedings of the 13th Conference on Language Resources and Evaluation (LREC 2022), Marseille, 20-25 June, 2022, França., Cited by: §2.
@@
- Natural language processing and machine learning for detection of respiratory illness by chest ct imaging and tracking of covid-19 pandemic in the united states. Radiology: Cardiothoracic Imaging 3 (1), pp. e200596. Cited by: §2.
@@
- Contributions to clinical information extraction in portuguese: corpora, named entity recognition, word embeddings. Master’s Thesis, Universidade de Coimbra (Portugal). Cited by: §2.
@@
- The role of data quality for reliable ai performance in medical applications. IEEE Reliability Magazine 1 (3), pp. 24–28. External Links: Document Cited by: §1.
@@
- IaraMed: a women’s healthcare chatbot for portuguese speakers. In Simpósio Brasileiro de Computação Aplicada à Saúde (SBCAS), pp. 931–942. Cited by: §2.
@@
- Leveraging online data to enhance medical knowledge in a small persian language model. arXiv preprint arXiv:2505.16000. Cited by: §1, §2.
@@
- Qcorp: an annotated classification corpus of chinese health questions. BMC medical informatics and decision making 18 (Suppl 1), pp. 16. Cited by: §3.3.
@@
- Applying deep matching networks to chinese medical question answering: a study and a dataset. BMC medical informatics and decision making 19 (Suppl 2), pp. 52. Cited by: §1, §2.
- Meddialog: two large-scale medical dialogue datasets. arXiv preprint arXiv:2004.03329. Cited by: §1.
@@
- Developing question-answering models in low-resource languages: a case study on turkish medical texts using transformer-based approaches. In 2024 8th International Artificial Intelligence and Data Processing Symposium (IDAP), Vol. , pp. 1–4. External Links: Document Cited by: §1, §2.
- A hybrid approach for named entity recognition in chinese electronic medical record. BMC medical informatics and decision making 19 (Suppl 2), pp. 64. Cited by: §2.
- Pubmedqa: a dataset for biomedical research question answering. arXiv preprint arXiv:1909.06146. Cited by: §1.
- MIMIC-iv, a freely accessible electronic health record dataset. Scientific data 10 (1), pp. 1. Cited by: §2.
- Robustness of ai-based prognostic and systems health management. Annual Reviews in Control 51, pp. 130–152. External Links: ISSN 1367-5788, Document, Link Cited by: §1.
- A review of applying large language models in healthcare. IEEE Access 13 (), pp. 6878–6892. External Links: Document Cited by: §1.
- BioRED: a rich biomedical relation extraction dataset. Briefings in Bioinformatics 23 (5), pp. bbac282. External Links: ISSN 1477-4054, Document, Link, https://academic.oup.com/bib/article-pdf/23/5/bbac282/45936115/bbac282.pdf Cited by: §2.
- Application of machine learning and word embeddings in the classification of cancer diagnosis using patient anamnesis. Ieee Access 8, pp. 106198–106213. Cited by: §2.
- MT-clinical bert: scaling clinical information extraction with multitask learning. Journal of the American Medical Informatics Association 28 (10), pp. 2108–2115. External Links: ISSN 1527-974X, Document, Link, https://academic.oup.com/jamia/article-pdf/28/10/2108/40408880/ocab126.pdf Cited by: §2.
- Overview and challenges of machine translation for contextually appropriate translations.” iscience 27 (10): 110878. DOI. Cited by: §1.
- Large language models in healthcare and medical domain: a review. informatics, 11, 57. Cited by: §1.
- Spbertqa: a two-stage question answering system based on sentence transformers for medical texts. In International Conference on Knowledge Science, Engineering and Management, pp. 371–382. Cited by: §1, §2.
- Medredqa for medical consumer question answering: dataset, tasks, and neural baselines. In Proceedings of the 13th International Joint Conference on Natural Language Processing and the 3rd Conference of the Asia-Pacific Chapter of the Association for Computational Linguistics (Volume 1: Long Papers), pp. 629–648. Cited by: §2.
- SemClinBr-a multi-institutional and multi-specialty semantically annotated corpus for portuguese clinical nlp tasks. Journal of Biomedical Semantics 13 (1), pp. 13. Cited by: §2.
- Adapting llms for the medical domain in portuguese: a study on fine-tuning and model evaluation. arXiv preprint arXiv:2410.00163. Cited by: §2.
- MedMCQA: a large-scale multi-subject multi-choice dataset for medical domain question answering. In Proceedings of the Conference on Health, Inference, and Learning, G. Flores, G. H. Chen, T. Pollard, J. C. Ho, and T. Naumann (Eds.), Proceedings of Machine Learning Research, Vol. 174, pp. 248–260. External Links: Link Cited by: §1.
- Toward structuring real-world data: deep learning for extracting oncology information from clinical text with patient-level supervision. Patterns 4 (4). Cited by: §1.
- The impact ofătraining data shortfalls onăsafety ofăai-based clinical decision support systems. In Computer Safety, Reliability, and Security, J. Guiochet, S. Tonetta, and F. Bitsch (Eds.), Cham, pp. 213–226. External Links: ISBN 978-3-031-40923-3 Cited by: §1.
- The metric-framework for assessing data quality for trustworthy ai in medicine: a systematic review. NPJ digital medicine 7 (1), pp. 203. Cited by: §1.
- Large language models encode clinical knowledge. arXiv preprint arXiv:2212.13138. Cited by: §1.
- Toward expert-level medical question answering with large language models. Nature Medicine 31 (3), pp. 943–950. Cited by: §2.
- Systematic evaluation of research progress on natural language processing in medicine over the past 20 years: bibliometric study on pubmed. Journal of medical Internet research 22 (1), pp. e16816. Cited by: §2.
- Deep learning in clinical natural language processing: a methodical review. Journal of the American Medical Informatics Association 27 (3), pp. 457–470. External Links: ISSN 1527-974X, Document, Link, https://academic.oup.com/jamia/article-pdf/27/3/457/34152802/ocz200.pdf Cited by: §2.
- Qwen3 technical report. arXiv preprint arXiv:2505.09388. Cited by: §5.1.
- Multi-scale attentive interaction networks for chinese medical question answer selection. IEEE Access 6 (), pp. 74061–74071. External Links: Document Cited by: §1.
- A unified framework of medical information annotation and extraction for chinese clinical text. Artificial Intelligence in Medicine 142, pp. 102573. External Links: ISSN 0933-3657, Document, Link Cited by: §1.
@@
Appendix A Question-Type Annotation Prompt
The following prompt (originally written in Brazilian Portuguese) was used to instruct gpt-oss-120b to classify each patient question into one of seven predefined semantic categories, as described in Section 3.3.